At 7am a train left the station heading directly north at 50 miles per hour. At 9am a bus left the same station heading directly south at 60 miles per hour. At what time will the bus & the train be exactly 650 miles apart?
I have tried to use the formula distance=rate x time but it doesn't seem to be correct.

Answer by stanbon(75887)

At 7am a train left the station heading directly north at 50 miles per hour.
At 9am a bus left the same station heading directly south at 60 miles per hour. At what time will the bus & the train be exactly 650 miles apart?
--------------------------------
North Train DATA:
rate = 50 mph ; time = x hrs; distance = 50x miles
-----------------------
South Train DATA:
rate = 60 mph ; time = x-2 hrs ; distance = 60(x-2) miles
----------
Equation:
distance + distance = 650 miles
50x + 60x-120 = 650
110x = 530
x = 4.82 hrs. (time for the north bound train)
----
Hour for north train = 7 + 4.82 hrs = 11.82
Note: 0.82hr = 0.82(60 min) is approximately 49 minutes
--------------
Ans: time = 11:49 AM
==========================
Cheers,
Stan H.